Each of two congruent sides of a isosceles triangle is 8in less than twice the base. The perimeter of the triangle is 74in What is the length of the base?
Answer by checkley71(8403)

X=BASE MEASUREMENT.
SIDE=2X-8.
X+(2X-8)+(2X-8)=74
5X-16=74
5X=74+16
5X=90
X=90/5
X=18 IS THE LENGTH OF THE BASE.
2*18-8=36-8=28 FOR THE SIDES
PROOF
18+2*28=18+56=74